Joe has a lot of juice. He has 10 liters of 80% apple juice. However, he wants to have less sugar so he adds water to his juice. How many liters of water should he add to make it 55% apple juice.
Answer by stanbon(75887)

Joe has a lot of juice. He has 10 liters of 80% apple juice. However, he wants to have less sugar so he adds water to his juice. How many liters of water should he add to make it 55% apple juice.
----------------
Equation::
juice + juice = juice
0.8*10 + 0*x = 0.55(10+x)
-------
80*10 + 0 = 55*10 + 55x
-----
800 - 550 = 55x
------
55x = 250
x = 50/11
x = 4 6/11 liters (amt. of water to add)
------------
Cheers,
Stan H.
============